In a class of 30 studnts, there are 17 girls and 13 boys. Five are A students, and three of these students are girls. If a student is chosen at random, what is the probability of choosing a girl and a A student?
Answer by stanbon(57307)

In a class of 30 studnts, there are 17 girls and 13 boys. Five are A students, and three of these students are girls. If a student is chosen at random, what is the probability of choosing a girl and a A student?
----------------------------
P(girl AND A-student) = P(girl)*P(A-student| girl)
= (17/30)*(3/17) = 3/30 = 1/10
=====================================
Cheers,
Stan H.